Planning Notebook

by Kenneth John Odle

v. 0.3.1 25 Octoer 2021

Contents

1 Introduction 1 1.1 Origins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 An Important Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 A Minor Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.4 Why Do People Have Trouble Keeping Track of Things? . . . . . . . . . . . . . . . . . . . . . . 1

2 Goals 2 3 Annual Goals 4 4 Monthly and Weekly Goals 5 5 Weekly Task List 7 6 The Year 9

Chapter 1

Introduction

I don’t like long introductions generally, so if this is to long for you, just read ”An Important Note” below and get on with life.

1.1 Origins

I created this for one main reason: I have a hard time getting and staying organized, especially with regard to financial matters. Part of this is no doubt due to my background (I did not grow up in an environment where people moved money around to benefit themselves) and undiagnosed ADHD (which is something I’m also looking into).

Also, I had a hard time making this and making it look nice with the tools I had available to me. (I like to do a lot of design things in my spare time and have spent countless hours questioning whether to go with a 10pt font size or a 10.5pt font size. The struggle is real.) Because I like learning new things, I decided to finally learn how to use L^{A}TEX. I have no idea what took me so long—the learning curve is not steep and it’s remarkably powerful. Because it’s basically a text file, I can compile it to a pdf file to share, but I can also share it via Git for those who want to create their own version.

1.2 An Important Note

You don’t have to wait until the beginning of a new year to start this. In fact, I suggest that you avoid New Year’s Day to make any major changes in your life. There is already so much going on, what with it being the height of the holiday season for most western religions. As well, people expect you to commit to changes. (And take delight in when you fail.)

Rather, start where you are, and make changes as gradually or as quickly as you feel comfortable with. As the cliché goes, there’s no time like the present.

1.3 A Minor Note

Like I mentioned above, I’ve generated this document using L^{A}TEX. As a result, this document is subject to all the limitations of L^{A}TEX, and more specifically, my limited (though growing) abilities with L^{A}TEX. If you want to correct any markup errors I’ve made (and I’m sure there are plenty), please contact me.

1.4 Why Do People Have Trouble Keeping Track of Things?

Goals

Let’s talk about goals.

Presumably you want to make a plan because you have some end in mind—something in your life that you want to change, see more of, or see less of. We’re going to work on a year-long plan to get you at least partway to those goals.

There are six parts to setting and achieving a goal, and if you studied journalism in high school or college, you’ll recognize them. Those six parts are what, who, when, where, why, and how. Let’s look at each one in a little bit more detailed.

What

This is the easy part. ”What” means what you actually want to achieve. This can be anything from ”lose ten pounds” to ”save �2,000 for a down payment on a new car”. Whatever you choose, it must be specific and measurable. In other words, it has to be something you can see. This is why most people who fail to achieve a goal: it’s not specific. You can easily tell if you’ve gained ten pounds just by getting on a scale. You can easily tell whether you’ve saved money by looking at your bank balance.

But if your goal is ”get more organized” there’s no way to measure ”organized”. The way out of this dilemma is to stop and think about what that’s going to look like.

Who

There are actually two “who”s here: Who can I rely on to help me achieve this goal, and who will benefit by my achieving this goal. With almost every goal, you’re going to be the main beneficiary. But often, there are other people in your life who will also benefit when you achieve this goal.

The other thing to consider is who you are going to tell about this goal. Yes, some goals may be very private, but making your goals public tends to help you achieve them.

How

The “how” describes the methods you are going to use to achieve this goal. And yes, it is methods, plural. If a goal is worth attaining, it’s probably going to require a few different techniques to achieve it. For example, if your goal is to lose weight, you will probably achieve it through a combination of diet and exercise. This will mean that you have to learn to cook healthier food and how to shop for healthier food, as well as learning how to exercise safely.

If your goal is to stop an unhealthy habit (such as smoking), then you are going to have to learn what to do in its place. This could also involve multiple techniques to achieve it.

Put a lot of thought into your “how.” This part of goal-setting is where a lot of your sub-goals will come from. But a lot of these will also end up as dead ends. You may have thought that a particular technique would work for you, but it doesn’t, so you’ll need to go with a different plan.

When

When do you want to achieve this goal by? One month, six months, 12 months? Or this a goal that is going to take more than a year? If so, where do you want to be at the end of the first 12 months?

You will also want to build some milestones into your timeline. A lot of people make a goal of exercising every day, and then when they skip a day, they figure that they will just exercise twice as much the next day. But when you put something off like that, it becomes easier and easier to just keep putting it off. The way to avoid this is by building milestones into your timeline.

For example, if your goal is to lose 30 pounds in the next nine months, then you should lose an average of just over three pounds per month. You should set some milestones so that after three months, you will have lost ten pounds, and after six months, you will have lost 20 pounds. This gives you some intermediate goals to aim for, but also provides some flexibility if you have a bad month.

In this planner, you will have space to list annual (i.e., 12-month) goals, and then space to break these down into smaller goals which you think of as monthly goals. Within the planner pages themselves, you will then have room to break down the monthly goals into goals for each week. This is where your “how” and “when” come together.

Where

Not many people give thought to where they are going to achieve their goals, as it is often not very important. But for some goals, location is everything. If you’re trying to eat healthier, for example, you may need to start going to a different grocery store, or take a different route home which avoids driving by so many fast food restaurants and their drive-throughs.

Why

This is where your motivation is going to come from. Why do want to achieve this goal? How will achieving it make your life better?

Chapter 3

Annual Goals

Okay, let’s make some goals for the next 365 days.

Chapter 4

Monthly and Weekly Goals

Goal #1

6

Weekly Task List

I find it helpful to keep track of what household chores need to be done. That way, I don’t spend an entire day each weekend just cleaning house and catching up on laundry, and I can plan ahead for needed supplies. If you have a busy household, you may find it handy to print out the following page, fill in the chores and assignments, and then laminate it and stick it on the front of your refrigerator. Everyone can then use a whiteboard marker to cross off their chore when it is done. You can then have a celebration on Sunday night when everything is done and you can wipe the entire sheet clean again, ready for the next week.

Weekly Tasks

Day  Task  Assigned To

Mo Tu We Th Fr Sa Su

Chapter 6

The Year

The rest of this planner consists of 52 weeks of planning pages.

Week 1 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 1 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 2 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 2 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 3 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 3 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 4 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 4 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 5 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 5 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 6 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 6 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 7 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 7 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 8 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 8 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 9 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 9 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 10 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 10 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 11 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 11 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 12 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 12 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 13 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 13 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 14 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 14 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 15 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 15 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 16 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 16 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 17 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 17 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 18 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 18 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 19 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 19 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 20 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 20 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 21 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 21 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 22 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 22 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 23 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 23 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 24 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 24 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 25 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 25 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 26 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 26 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 27 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 27 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 28 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 28 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 29 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 29 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 30 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 30 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 31 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 31 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 32 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 32 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 33 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 33 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 34 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 34 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 35 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 35 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 36 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 36 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 37 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 37 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 38 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 38 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 39 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 39 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 40 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 40 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 41 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 41 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 42 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 42 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 43 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 43 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 44 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 44 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 45 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 45 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 46 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 46 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 47 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 47 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 48 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 48 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 49 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 49 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 50 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 50 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 51 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 51 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For:

Week 52 Month and Year: Date Tasks Today/Appointments Financial

Communications

Mo Tu We Th Fr Sa Su

Goals for this Year Goals for this Month

Hoping For:

Week 52 Look For: Health Tasks This Week

Notes

Goals for this Week

Grateful For: